OpenStudy (anonymous):

Solve for x: 2|x − 3| + 1 = 7

OpenStudy (zzr0ck3r):

first get it down to |x-3| = something and tell me what you have

OpenStudy (anonymous):

@zzr0ck3r, since your so smart, can you help @jinan

OpenStudy (anonymous):

x-3=3

OpenStudy (zzr0ck3r):

|x-3|=3

OpenStudy (anonymous):

would it be x-3=3 and x-3=-3?

OpenStudy (zzr0ck3r):

ok now when we have something of the form |a|=b we will have two equations a=b and a=-b so remove the absolute value signs and do that

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

solve for x in both equations

OpenStudy (anonymous):

okay thanks!

OpenStudy (zzr0ck3r):

@eric21 np

OpenStudy (anonymous):

ok, can you see if you can figure it out?

OpenStudy (anonymous):

hhhhhhh 4sodapop who told you that im smart

OpenStudy (anonymous):

@jinan? i was saying that @zzr0ck3r seems smart and seeing if they could help you.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

you right i need help @zzr0ck3r